OpenStudy (anonymous):

\[\log _{6}1/\sqrt{6}\] help me ???

OpenStudy (anonymous):

log 1 - log sqrt 6

OpenStudy (anonymous):

\[\log _{1}-\log ^{6}\] like this ???

OpenStudy (anonymous):

no, log (a/b) = log a =log b so\[\log_{6}\frac{1}{\sqrt{6}} = \log_6 1 - \log_{6} \sqrt{6}\]

OpenStudy (anonymous):

whoops sorry, there is an easier way to do this.

OpenStudy (cruffo):

Recall that \[\sqrt{6} = 6^{1/2}\] so you can simplify more...

OpenStudy (anonymous):

\[\log_{6}\frac{1}{\sqrt{6}} = \log_{6} 6^{\frac{-1}{2}}\]

OpenStudy (anonymous):

\[= \frac{-1}{2}\log_66 = -1/2\]

OpenStudy (anonymous):

so the last comment is the answer??? or just the -1/2

OpenStudy (anonymous):

\[\log_nn=1\]